Consider a bank with the following income statement: It has $100 in loans with an

interest rate of 5 percent; $50 in security holdings, paying 10 percent; noninterest

income of $10; $100 in savings accounts that have an interest rate of 2.5 percent; and

other expenses of $15. How much profit does this bank make?

A) –$5

B) $2.50

C) $35

D) $20

Which of the following is a “core deposit”?

A) checking deposits

B) savings deposits

C) small-time deposits

D) All of the answers are correct.

The main reasons why the banking business has evolved include:

A) competition from securities markets.

B) deregulation.

C) financial innovation.

D) All of the answers are correct.

Market risk arises from:

A) business cycle fluctuations.

B) asset price fluctuations.

C) oil price changes.

D) interest rate volatility.
Answer #1


Question 1

Bank has $100 in loans. It charges an interest rate of 5 percent.

Amount of interest charges on loans = $100 * 0.05 = $5

The amount of interest charged on loans is $5.

Bank has $50 in security holding. Securities earn 10% interest.

Return on securities = $50 * 0.10 = $5

The bank earn $5 as interest from securities.

Non interest income is $10.

Bank has $100 as saving account deposit. It pay interest of 2.5%.

Interest paid on savings account = $100 * 0.025 = $2.5

The interest paid on savings account is $2.5

Other expenses is $15.

Calculate the profit of bank -

Profit = Aamount of interest charged on loans + Return on securities + Non interest income - Interest paid on savings account - other expenses

Profit = $5 + $5 + $10 - $2.5 - $15 = $2.50

The profit of this bank is $2.50

Hence, the correct answer is the option (B)

Note -: As per the Chegg Answering Policy, when more than 1 question is posted then, in that case, only 1st posted question is answered with complete explanation.